Start-up Observing Night 4 - 20160911 UT

Observers: JHill, DMiller, BRothberg, AVaz, PHinz (Tucson)
Telescope Operator: GBechetti (LBTO)
AO Support: DMiller
Telescope Support: MWagner (LBTO)
Software Support: KSummers (Tucson)
Instrument Support: JMorris (LBTO)
Engineering Support:

Summary

Beautiful mostly clear sky with 0.5 arcsec seeing for much of the night. Completed AGw1 (and AGw2) probe alignment - both were already good. Completed AGw1 and AGw2 M3 adjustment for pupil wobble. Collected field aberration data on both sides. Working on W-unit transform for AGW1.

Preparations

Switched forward to Build 2016B this afternoon with AOS sorted out.

TO DO following this night: (for JMH)
Restart OSS to get updated M3 positions in positions.conf, and clear the global offsets in PSF. DONE
Install the binodal corrections in the collimation models. DONE
Update the pointing models to include both of the above.

Details

Open

LBTI in twilight

01:30 Using bright twilight to recover LBTI pointing and pupil positions on WFS.

Nice twilight pupils on SX, but can't find Vega after relatively large spiralling.

Pupils on DX look vignetted with a circular edge, but easily found Vega (in focus on WFS) at a similar pointing to LUCI. DX IE=-14 CA=-245 Perhaps the DX dichroic didn't get remounted at the correct angle?

AGw1 and AGw2 Alignment of Probe to Rotator Center

02:13 Authorize back to LUCI/LUCI

ACQUIRE

02:29 ACTIVE preset to BS9177 - but SX pointing changed ~10 arcec. ?? Seeing is ~0.5 arcsec.

ACQUIRE then Home M2s and M3s.

Resend ACTIVE preset.

See BGTechProcedures for instructions.

data in /lbt/data/share/tcs/telescope.20160911/

On each side: Stop guiding
Move rotator to 360.
ptincrement 0 CA -5
gcsclient readGuideCam left -e 1000 (test fails)
gcsclient readGuideCam left -e 1000 (OK)
gcsclient readGuideCam left -e 90000
Move rotator to 0

SX Hotspot is 241,248.5 in absolute CCD pixels.

SX guide 00002 center is 115,105 binned, 230,210 unbinned, 245,249 detsec corrected.

DX Hotspot is 275,243 in absolute CCD pixels.

DX guide 00002 center is 127.2.115.4 binned, 254.4,230.8 unbinned, 271.4,245.8 detsec corrected.

These are both within the uncertainties of "perfectly centered", so we declare victory and move on to the next task.

AGw1/2 Pupil Wobble

resend ACTIVE preset to collimate

Data is in * /home/lbto/data/20160911/ * Both sides were orignally in that same directory, but JMH moved the wobble files into sub-directories SX_wobble and DX_wobble to make analysis easier.

first on LFBG, then start in parallel on RFBG using another IDL session

iif_register, LFBG

press "stop WFSing"

pupil_collect, retval, EXP=10.0 (12th mag star)

03:23 RFBG failed to recover guide star on offset. Why?

03:27:20 Telescope off source cancels preset - rotator failed? Restart LFBG rotator. Resend preset. Same problem happened to RFBG rotator, but it won't recover.

03:38 pupil_collect, retval, EXP=10.0 (on LFBG), but LFBG rotator failed again.

03:44 Restart mcspu. Then power cycle the rotator rack in treehouse. Then restart mcspu again.

04:09 ACQUIRE M5_0738

04:15 Resend preset to BS9177

04:18 pupil_collect, retval, EXP=10.0 (on LFBG, and then RFBG)

Absolute SX Tip=645 Tilt=90 Selector=29.798 Absolute DX Tip=-900 Tilt=100 Selector=-25.721 (M3 Global Tip-Tilt at 0,0)

Left missed the star on rotator offset by 1 arcsec, and right by ~3.5 arcsec (because DX has an offset on the default preset).

Block sending centroids to PCS
Move M3 Tip/Tilt in Global Offset Increments of 50 arsec
Recover pointing with IE/CA from Paddle
Resume sending centroids to PCS
stop WFSing
Do another round of pupil_collect.

Doug sends DX preset without the hotspot. (irc_sethotspot) Seeing is 0.42-0.55 arcsec.

RFBG rotator is at 429 degrees.

04:40 M3 Global Tip=+200 arcsec on both sides

05:00 M3 Global SX Tip=-400 DX Tip -200

05:15 M3 Global SX Tip=-400 Tilt=-200 DX Tip=-100

05:25 M3 Global SX Tip=-500 Tilt=-100 DX Tip=-100 Tilt=-100 (finished with SX)

05:39 Resend preset to unwrap rotators.

05:45 M3 Global SX Tip=-500 Tilt=-100 DX Tip=-100 Tilt=100

Note that the DX wobble plot converges to a line extended in the X-direction, as opposed to the very small circle that you expect. This is the same behaviour we saw last time (Oct 2015?) on DX.

Decide to Apply Global Offsets * SX Tip=-500 Tilt=-100 DX Tip=-100 Tilt=0 *
(Absolute SX Tip=+145 Tilt=-10 Selector=29.798 DX Tip=-1000 Tilt=+100 Selector -25.720)
JMH applied these to LFBG and RFBG in positions.conf, but we have not yet restarted OSS.

06:02 ACQUIRE ---> SX IE=-3 CA=-47 DX IE=+22 CA=-242

After the fact analysis 20201005
SX
filename              Xavg     Yavg     Radius   pts
pupil_05:28:11.wobble  383.315  214.935  1.61059 4
pupil_05:16:07.wobble  382.560  215.028  2.93950 4
pupil_05:01:38.wobble  382.688  216.325  4.66193 4
pupil_04:40:41.wobble  379.843  215.400 17.15972 4
pupil_04:18:31.wobble  380.862  216.153 12.75017 4
Screenshot at 2020-09-03 22-20-17.png

DX
filename              Xavg     Yavg     Radius   pts
pupil_05:48:34.wobble  377.035  268.665  9.66330 4
pupil_05:29:49.wobble  376.658  271.480 10.27528 4
pupil_05:16:20.wobble  378.140  270.142  8.56166 4
pupil_05:05:32.wobble  377.543  268.982  9.27712 4
pupil_04:44:28.wobble  377.683  270.183 14.09598 4
pupil_04:18:45.wobble  377.548  270.433  8.23036 4
Screenshot at 2020-10-05 22-21-58.png

Big W Transform on LUCI1 AGW1 (part 1)

see transform details in IT 5435

06:07 Barry sends LUCI1 preset to M29, and telescope moves with ACQUIRE preset running on right. (LUCI is sending the CLEARPRESET flag - see below.) Missed left gs off left edge of acq field. Bump SX IE -5 with paddle. Resend preset - OK.

Sun Sep 11 06:13:19.413 2016 left RunAO alert for LUCIFER: AcquireRef{FLAO}[45] - RETRY: source_acq error: mag -11.678 - expected: 13.698 difference: 25.376) (-20002) WFSARB_STAR_WRONGFLUX
Sun Sep 11 06:13:19.424 2016 left RunAO command failed from LUCIFER, type NORMAL, leftRePoint 0, rightRePoint 0

Perhaps no star on CCD47? Wrong X-direction??

06:31 Doug sends Binocular preset to OC_M29 (aka HD229234)

* There is likely more information in Doug's AO wiki log. We are running parallel logs tonight as we are sometimes doing different tasks. *

6.55 to 6.70 UT Very strange guiding oscillation on SX only. Cause is unknown.

07:13 AO preset using target as off-axis AO reference.

DX FLAO Bayside patrol limits are a rectangle from -38.0,-30.0 to +74.0,50.0 mm.

08:10 ACQUIRE preset

08:40 LUCI preset to M34

Sun Sep 11 08:40:21.162 2016 both BinocularControl command complete for LUCIFER with flag CLEARPRESET

Didn't find ref star, but guide star was a wide binary.

Doug sees a band of vignetting at radius 44 arcsec in the AO patrol field.

08:45 ACQUIRE

09:14 Problems with the LUCI transform script (not resending the preset for each reference star), so we switch to taking field aberration data while working on an improved script.

Field Aberration Data on Both Sides BS9107

09:16 ACQUIRE SYNC BS9107

09:18 field_collect, filename, LIMIT=200.0 (BS9107 new field) Did this on both sides in parallel asynchronously.

09:2x Restart field_collect to avoid race condition when they are running in same directory. Seeing is about 0.65 arcsec at EL=70 deg.

09:45 Reflatten DX AdSec and Clear Active Optics on M1 to avoid collimation anomalies that we think are being caused by wrinkles in the shell. That made the wfs spots look much better, but didn't really eliminate the ~1.2 microns of focus change when we move back on-axis.

Doug suggests SX Tip=+20 Tilt=0 dIE=-20, DX Tip=-20 Tilt=-20 dIE=+20 dCA=+20. See the AO wiki log for details.

Big W Transform on LUCI1 AGW1 (part 2)

10:34 Back to W-Unit Transformation with a new LUCI1 script

10:34 ACE_ACTIVE preset to M34

The CLEARPRESET seems to have corrupted the previously send ACTIVE preset on right by moving the telescope to a new field. That cause PSF Wendy to become distressed.

Sun Sep 11 10:31:45.749 2016 right Preset command started for LUCIFER mode ACTIVE with guidestars, hotspot
Sun Sep 11 10:31:48.052 2016 right Preset command wait for elevation started for LUCIFER
Sun Sep 11 10:31:49.563 2016 right Preset command wait for elevation complete for LUCIFER
Sun Sep 11 10:31:57.665 2016 right Preset command wait for telescope started for LUCIFER
Sun Sep 11 10:31:57.679 2016 right Preset command wait for telescope complete for LUCIFER
Sun Sep 11 10:32:40.311 2016 right Preset command complete for LUCIFER mode ACTIVE with guidestars, hotspot, Sound=presetComplete.wav
Sun Sep 11 10:33:52.459 2016 right Preset command started for LUCIFER mode ACTIVE with guidestars, hotspot
Sun Sep 11 10:33:56.088 2016 right Preset command wait for elevation started for LUCIFER
Sun Sep 11 10:33:57.604 2016 right Preset command wait for elevation complete for LUCIFER
Sun Sep 11 10:34:05.698 2016 right Preset command wait for telescope started for LUCIFER
Sun Sep 11 10:34:29.325 2016 both BinocularControl command started for LUCIFER with flag CLEARPRESET
Sun Sep 11 10:34:29.493 2016 both BinocularControl command complete for LUCIFER with flag CLEARPRESET
Sun Sep 11 10:34:29.609 2016 left Preset command started for LUCIFER mode ADAPTIVEACE_ACTIVE with guidestars, offset, AO reference
Sun Sep 11 10:34:31.874 2016 left Preset command wait for elevation started for LUCIFER
Sun Sep 11 10:34:44.894 2016 left Preset command wait for elevation complete for LUCIFER
Sun Sep 11 10:34:57.420 2016 left Preset command wait for telescope started for LUCIFER
Sun Sep 11 10:36:02.439 2016 left Preset command wait for telescope complete for LUCIFER
Sun Sep 11 10:36:02.721 2016 right Preset command wait for telescope complete for LUCIFER
Sun Sep 11 10:36:42.788 2016 left Preset command complete for LUCIFER mode ADAPTIVEACE_ACTIVE with guidestars, offset, AO reference, Sound=presetComplete.wav
Sun Sep 11 10:37:11.163 2016 right Preset alert for LUCIFER mode ACTIVE: Stop{FLAO}[115]
Sun Sep 11 10:37:11.185 2016 right Preset alert for LUCIFER mode ACTIVE: guiding was not stable within 6 cycles or lost GS
Sun Sep 11 10:37:11.200 2016 right Preset command failed for LUCIFER mode ACTIVE with guidestars, hotspot, Sound=presetFailed.wav
Sun Sep 11 10:37:14.975 2016 left RunAO command started from LUCIFER, type NORMAL, leftRePoint 0, rightRePoint 0
Sun Sep 11 10:38:35.089 2016 left RunAO alert for LUCIFER: AcquireRef{FLAO}[130]
Sun Sep 11 10:38:35.098 2016 left RunAO command failed from LUCIFER, type NORMAL, leftRePoint 0, rightRePoint 0

11:00 First transform image is actually collected by LUCI1!

Starting exposure is #19 through #25.

Close

12:20 Close for impending dawn. Still spectroscopically clear with humidity up to 80%.

12:26 Stop and restart OSS to read the new positions.conf. Removed equivalent global offsets from PSF M3.

-- %USERSIG{JohnHill - 2016-09-11}%
Topic attachments
I Attachment Action Size Date Who Comment
Screenshot at 2020-09-03 22-20-17.pngpng Screenshot at 2020-09-03 22-20-17.png manage 34 K 05 Oct 2020 - 22:58 JohnHill SX M3 Pupil Wobble Plot
Screenshot at 2020-10-05 22-21-58.pngpng Screenshot at 2020-10-05 22-21-58.png manage 37 K 05 Oct 2020 - 22:26 JohnHill DX M3 Pupil Wobble Plot
Topic revision: r5 - 05 Oct 2020, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback